OpenStudy (anonymous):

what is the pressure of a 50.0L tank that contains 4.04kg of oxygen, at 23 degrees celsius?

OpenStudy (anonymous):

we are going to use PV = nRT eq. P(pressure) = ?; V = 50.0L; R = 0.082(if P is in atm), or 8.314(if P is in J) T=23C + 273 = 296K n(O2) = m(O2)/Mr(O2) = 4.04/32~0.13 P * 50 = 0.13 * 0.082(or 8,314) * 296 P = 0.063atm(or 6.4J)